Write the equation of the line passing through each of the given points with the indicated
slope. Give your results in slope-intercept form, where possible.
(0,5),m=-3/5

Answer by tutorcecilia(2152)

(0,5),m=-3/5
(y-y1)=m(x-x1) [begin with the point-slope form]
(y-5)=(-3/5)(x-0) [simplify]
(y-5)=-3/5x-(-3/5)(0)
y-5=-3/5x [solve for y]
y=-3/5x+5
.
check by plugging point (0,5) back into the equation]
y=-3/5x+5
5=-3/5(0)+5
5=5 [checks out]